Felix Drove 800 miles in X amount of hours on Monday.
A) Write a rational expression for his average speed
B) On tuesday he drove for 6 hours at the same average speed.
Write a rational expression for his distance.... Please i need help

Answer by jsmallt9(3759)

A) Average speed = (total distance)/(total time)
So Felix's average speed is 800/x

B) Distance = (average speed) * (total time)
So Felix's distance on Tuesday = (800/x)*6 = 4800/x